# Partial pressure and susceptibilities of charmed sector in the van der Waals hadron resonance gas model

### v1

@@ II.4 Partial pressure and Charm susceptibilities
$P_{B}^{C}$ $=\frac{1}{2}\left(3\chi_{22}^{BC}-\chi_{13}^{BC}\right)\,$, (11)
$P_{M}^{C}$ $=\chi_{4}^{C}+3\chi_{22}^{BC}-4\chi_{13}^{BC}\,.$
@@
$P_{M}^{C,S=1}$ $=\chi_{13}^{SC}-\chi_{112}^{BSC}\,$, (12)
$P_{B}^{C,S=1}$ $=\chi_{13}^{SC}-\chi_{22}^{SC}-3\chi_{112}^{BSC}\,$,
$P_{B}^{C,S=2}$ $=\frac{1}{2}\left(2\chi_{112}^{BSC}+\chi_{22}^{SC}-\chi_{13}^{SC}\right)\,.$
@@
$T(\mu_{B})=q_{1}-q_{2}\mu_{B}^{2}-q_{3}\mu_{B}^{4},$
$\mu_{B}(\sqrt{s_{\text{NN}}})=\frac{q_{4}}{1+q_{5}\sqrt{s_{\text{NN}}}},$
where, $q_{1}=0.166~\text{GeV}$, $q_{2}=0.139~\text{GeV}^{-1}$, $q_{3}=0.053~\text{GeV}^{-3}$, $q_{4}=1.308~\text{GeV}$, and $q_{5}=0.273~\text{GeV}^{-1}$. These parameters are obtained using freeze-out criteria based on the ideal HRG model. Although similar estimations have been made using the EVHRG and VDWHRG model, the resulting parameter variations are negligible [42, 43].
@@ III Results
For this work, we have used all the charmed hadrons in the PDG [44], as well as the undiscovered charmed states from the quark model [45, 46]. In the left panel of Fig. 1, we present the partial pressure of charmed baryons, estimated using the quantity $(3\chi_{22}^{BC}-\chi_{13}^{BC})/2$. While the HRG model, including contributions from undiscovered states, overestimates this quantity and deviates from lQCD data for $T\gtrsim 160$ MeV, the VDWHRG model provides an excellent description up to $T\sim 180$ MeV. This suggests that the inclusion of attractive and repulsive interactions between hadrons, as implemented in the VDWHRG model, is crucial for describing the thermodynamics of charmed baryons near the crossover region. Similarly, in the right panel of Fig. 1, we use Eq. 11 to estimate the partial pressure of charmed mesons as a function of temperature. We observe an increasing trend in the lattice QCD estimations throughout the entire temperature range, which is well described by all the HRG models. However, we still notice that incorporating interactions among the hadrons decreases the partial pressure of the charmed mesons. Moreover, in the left panel of Fig. 2, we plot the partial pressure of the charmed baryon as a function of temperature and baryo-chemical potential. We observe that along the $\mu_{\text{B}}$ axis, the partial pressure of charmed baryon increases at low temperature, but drops slightly for higher $\mu_{\rm{B}}$ at higher temperatures. Additionally, in the right panel, we observe a monotonically increasing trend for the partial pressure of charmed mesons with increasing temperature. In comparison, with increasing baryo-chemical potential, the partial pressure of charmed mesons is unaffected.
@@
Now, in the left panel of Fig. 3, we study the partial pressure of charmed mesons with strangeness one. The IHRG, EVHRG, and VDWHRG models describe the lQCD data reasonably well, though the EVHRG and VDWHRG models provide a better match at high temperatures. The relatively good performance of the models for these states may be attributed to the fact that most of the expected charm-strange mesons (e.g., $D_{s}$ mesons) have been experimentally observed, allowing for a more complete implementation in the hadron resonance gas framework. In the middle panel of Fig. 3, we plot the partial pressure of charmed baryons with strangeness one. In this case, all the HRG models match the lQCD data within the large error bars. Although the EVHRG seems to show a reasonable agreement with the lattice QCD data at higher temperatures. Similarly, the right panel of Fig. 3 shows charmed baryons with strangeness two, and both the IHRG and VDWHRG models give similar values, which underpredict the lQCD estimations. In addition, the EVHRG shows even worse agreement with lQCD. This might be due to the fact that a larger number of charm-strange baryons, with $|S|=2$, are still undiscovered and not predicted by the quark models yet. The similar values of IHRG and VDWHRG point to the fact that the effect of VDW is smaller as the number density of charmed-baryons with strangeness two is very low, yet the attractive interactions counter the repulsive interactions reasonably. Future experimental efforts at facilities like LHCb and the upcoming Electron-Ion Collider (EIC) could help constrain these states better.
In Fig. 4 (left panel), we plot the ratio of partial pressure from charmed mesons with strangeness one to total charmed pressure. For the hadron gas case, we have used total charm pressure as the added pressure from charmed mesons and charmed baryons. However, for the lQCD, they have also added a charm quark pressure, assuming a mixed state near the transition temperature. Due to this contribution from charm quark pressure, the lQCD trend goes down after $T\sim$ 160 MeV. The VDWHRG model explains the lQCD data up to $T\sim$ 160 MeV, but then shows an upward trend after that. This mismatch comes from the neglected charm quark pressure, which seems to dominate after the transition temperature.
Figure 6: Charm susceptibilities ratios, $\chi_{2}^{C}/\chi_{1}^{C}$ (left) and $\chi_{4}^{C}/\chi_{3}^{C}$ (right), as a function of $\sqrt{s_{\rm NN}}$.
Figure 7: Charm susceptibilities ratios, $\chi_{3}^{C}/\chi_{1}^{C}$ (left), $\chi_{3}^{C}/\chi_{2}^{C}$ (middle), and $\chi_{4}^{C}/\chi_{2}^{C}$ (right), as a function of $\sqrt{s_{\rm NN}}$.
@@
On the other hand, the charm susceptibilities and their ratios as functions of collision energy reveal a multi-faceted picture of QCD matter across different phases. Much like in the baryon sector, in Fig. 6 the increasing trends of $\chi_{2}^{C}/\chi_{1}^{C}$ and $\chi_{4}^{C}/\chi_{3}^{C}$ with $\sqrt{s_{\rm NN}}$ demonstrate the gradual emergence of correlated charm production mechanisms, transitioning from independent fragmentation at low energies to medium-dominated processes in the QGP phase at high energies. The consistent model hierarchy (VDWHRG << IHRG << EVHRG) highlights how van der Waals interactions most effectively capture the partial thermalization of charm quarks, with attractive forces suppressing fluctuations closer to lQCD predictions. This is complemented by the constancy of $\chi_{3}^{C}/\chi_{1}^{C}$ and $\chi_{4}^{C}/\chi_{2}^{C}$ at unity across all energies as shown in the left and right panels of Fig. 7, which serve as a thermodynamic fingerprint of charm’s unique properties, showing that its large mass preserves the charm symmetry and Gaussian statistics despite the phase transition. The decreasing trend of $\chi_{3}^{C}/\chi_{2}^{C}$ as shown in the middle panel of Fig. 7 provides additional evidence of the hadron-QGP crossover, with its saturation at high energies reflecting the decoupling of charm from baryon number in the $\mu_{B}\to 0$ regime. This comprehensive view positions charm fluctuations as both a sensitive probe of the QCD phase diagram and a stringent test for theoretical models across the entire collision energy spectrum.
@@
The systematic evolution of these ratios with collision energy offers a powerful tool for mapping the QCD phase diagram. Future measurements with improved precision could further constrain the location of the critical point and clarify the temperature dependence of charm hadronization. In addition to this, one key feature that could substantially change these ratios is the charm chemical potential. At RHIC energies, the charm chemical potential ($\mu_{c}$) can be finite due to the nature of charm quark production and conservation in heavy-ion collisions. Charm quarks are predominantly produced in initial hard scatterings rather than through thermal excitation, and their number remains approximately conserved throughout the evolution of the fireball. In thermal models like the HRG, a finite $\mu_{c}$ can be introduced to enforce this conservation, especially when using a grand canonical ensemble. Moreover, local imbalances between charm and anti-charm quarks can develop due to event-by-event fluctuations or limited charm diffusion, further justifying a nonzero $\mu_{c}$ [47]. In thermal fits to charm hadron yields (such as $D^{0}$, $\Lambda_{c}$, or $J/\psi$), incorporating a finite charm chemical potential is necessary to correctly reproduce observed yields under the constraint of fixed net charm. Thus, a thorough look into charm chemical potential and subsequent charm susceptibility estimation is very timely and necessary.
@@ IV Summary
In this work, we study the thermodynamic properties and charm sector fluctuations of QCD matter using the van der Waals Hadron Resonance Gas (VDWHRG) model. By incorporating both attractive and repulsive interactions between hadrons, the VDWHRG model significantly improves agreement with lattice QCD (lQCD) data, especially in the temperature range up to 180 MeV. We show that the VDWHRG model, by incorporating both attractive and repulsive hadronic interactions, provides an improved description of charm susceptibilities and partial pressures. Our study includes the estimation of partial pressures of charmed mesons and baryons, both with and without strangeness, and compares them to available lQCD results. While VDWHRG captures the qualitative trends well, especially below $T\sim 160$ MeV, discrepancies at higher temperatures highlight the possible role of deconfined charm degrees of freedom not accounted for in hadron-based models. Additionally, we extend our analysis to baryon-rich regimes, offering insights into charm susceptibilities in regions not easily accessible by lQCD.
Furthermore, we study the charm susceptibilities and their ratio as a function of collision energy, which highlights their sensitivity to the underlying QCD phase structure. The observed trend for $\chi_{C}^{2}/\chi_{C}^{1}$ and $\chi_{C}^{4}/\chi_{C}^{3}$, shows the gradual increase in charm production with an increase in collision energy. Moreover, a saturation behavior in $\chi_{C}^{3}/\chi_{C}^{2}$ highlights the decoupling of the charm and baryon number at higher center-of-mass collision energy. To summarize, our results emphasize the importance of including both attractive and repulsive interactions among the hadrons for a more complete study of charm thermodynamics within the hadronic sector. With the upcoming facilities, such as the Future Circular Collider and EIC, providing robust theoretical constraints on the charm sector production will be crucial to the experimental efforts.

### v2

@@ II.4 Partial pressure and Charm susceptibilities
$P_{B}^{C}$ $=\frac{1}{2}\left(3\chi_{22}^{BC}-\chi_{13}^{BC}\right)\,$, (11a)
$P_{M}^{C}$ $=\chi_{4}^{C}+3\chi_{22}^{BC}-4\chi_{13}^{BC}\,.$ (11b)
@@
$P_{M}^{C,S=1}$ $=\chi_{13}^{SC}-\chi_{112}^{BSC}\,$, (12a)
$P_{B}^{C,S=1}$ $=\chi_{13}^{SC}-\chi_{22}^{SC}-3\chi_{112}^{BSC}\,$, (12b)
$P_{B}^{C,S=2}$ $=\frac{1}{2}\!\left(2\chi_{112}^{BSC}+\chi_{22}^{SC}-\chi_{13}^{SC}\right)\,.$ (12c)
@@
$T(\mu_{B})=q_{1}-q_{2}\mu_{B}^{2}-q_{3}\mu_{B}^{4},$ (14)
$\mu_{B}(\sqrt{s_{\text{NN}}})=\frac{q_{4}}{1+q_{5}\sqrt{s_{\text{NN}}}},$ (15)
where, $q_{1}=0.166~\text{GeV}$, $q_{2}=0.139~\text{GeV}^{-1}$, $q_{3}=0.053~\text{GeV}^{-3}$, $q_{4}=1.308~\text{GeV}$, and $q_{5}=0.273~\text{GeV}^{-1}$. These parameters are obtained using freeze-out criteria based on the ideal HRG model. Although similar estimations have been made using the EVHRG and VDWHRG model, the resulting parameter variations are negligible [42, 43, 44]. Moreover, one should take note of the fact that this freeze-out parametrization determined from the yields of light-flavor hadrons. We work with the assumption that the charm hadrons decouple from the medium at the same temperature as light hadrons. For a comparison on equal footing, we employ the experimental $p_{\rm{T}}$ and $\eta$ cuts in the estimation of charm susceptibilities. The pressure integral in the hadronic models modifies as [45],
$P_{i}^{id}=\pm\frac{Tg_{i}}{4\pi^{2}}$ $\int_{-\eta^{min}}^{\eta^{max}}d\eta\int_{p_{\rm T}^{min}}^{p_{\rm T}^{max}}dp_{\rm T}~p_{\rm T,i}^{2}\cosh{\eta}$ (16)
$\times{\rm ln}\!\left(1\pm e^{-(E^{i}(p_{\rm{T}},\eta)-\mu_{i})/T}\right),$
here, $E^{i}(p_{\rm{T}},\eta)$ is the energy of the $i^{th}$ hadron given as, $E_{i}(p_{T},\eta)=\sqrt{p_{\rm T,i}^{2}\cosh^{2}\eta+m_{i}^{2}}$. Moreover, the chemical potentials are constrained by imposing $n_{Q}/n_{B}=0.4$, where $n_{Q}$ and $n_{B}$ are the net electric charge and net baryon densities, along with strangeness and charm neutrality. Furthermore, we incorporate the suppression of fluctuations arising from exact global conservation of the conserved charges, we correct the susceptibilities using the subensemble acceptance method [46]. In this framework, only a fraction of the total system contributes to the measured susceptibilities, leading to a suppression in the higher-order cumulants as compared to the grand-canonical values. The acceptance factor, $\alpha$, is given as $\alpha=N_{\rm{ch}}(\Delta p_{\rm{T}}\Delta y)/N_{\rm{ch}}(\infty)$ [46, 47]. It modifies the susceptibilities as,
$\frac{\kappa_{2}}{\kappa_{1}}$ $=\beta\,\frac{\chi_{2}}{\chi_{1}},$ (17)
$\frac{\kappa_{3}}{\kappa_{2}}$ $=\frac{(1-2\alpha)\,\chi_{3}}{\chi_{2}},$ (18)
$\frac{\kappa_{3}}{\kappa_{1}}$ $=\frac{\beta(1-2\alpha)\,\chi_{3}}{\chi_{1}},$ (19)
$\frac{\kappa_{4}}{\kappa_{3}}$ $=\frac{(1-3\alpha\beta)\,\chi_{4}-3\alpha\beta\,\dfrac{\chi_{3}^{2}}{\chi_{2}}}{(1-2\alpha)\,\chi_{3}},$ (20)
$\frac{\kappa_{4}}{\kappa_{2}}$ $=\frac{(1-3\alpha\beta)\,\chi_{4}-3\alpha\beta\,\dfrac{\chi_{3}^{2}}{\chi_{2}}}{\chi_{2}}.$ (21)
where, $\beta=(1-\alpha)$.
@@ III Results
For this work, we have used all the charmed hadrons in the PDG [48], as well as the undiscovered charmed states from the quark model [49, 50]. In the left panel of Fig. 1, we present the partial pressure of charmed baryons, estimated using the quantity $(3\chi_{22}^{BC}-\chi_{13}^{BC})/2$. While the HRG model, including contributions from undiscovered states, overestimates this quantity and deviates from lQCD data for $T\gtrsim 160$ MeV, the VDWHRG model provides an excellent description up to $T\sim 180$ MeV. This suggests that the inclusion of attractive and repulsive interactions between hadrons, as implemented in the VDWHRG model, is crucial for describing the thermodynamics of charmed baryons near the crossover region. Similarly, in the right panel of Fig. 1, we use Eq. 11b to estimate the partial pressure of charmed mesons as a function of temperature. We observe an increasing trend in the lattice QCD estimations throughout the entire temperature range, which is well described by all the HRG models. However, we still notice that incorporating interactions among the hadrons decreases the partial pressure of the charmed mesons. Moreover, in the left panel of Fig. 2, we plot the partial pressure of the charmed baryon as a function of temperature and baryo-chemical potential. We observe that along the $\mu_{\text{B}}$ axis, the partial pressure of charmed baryon increases at low temperature, but drops slightly for higher $\mu_{\rm{B}}$ at higher temperatures. Additionally, in the right panel, we observe a monotonically increasing trend for the partial pressure of charmed mesons with increasing temperature. In comparison, with increasing baryo-chemical potential, the partial pressure of charmed mesons is unaffected.
@@
Now, in the left panel of Fig. 3, we study the partial pressure of charmed mesons with strangeness one. The IHRG, EVHRG, and VDWHRG models describe the lQCD data reasonably well, though the EVHRG and VDWHRG models provide a better match at high temperatures. The relatively good performance of the models for these states may be attributed to the fact that most of the expected charm-strange mesons (e.g., $D_{s}$ mesons) have been experimentally observed, allowing for a more complete implementation in the hadron resonance gas framework. In the middle panel of Fig. 3, we plot the partial pressure of charmed baryons with strangeness one. In this case, all the HRG models match the lQCD data within the large error bars. Although the EVHRG seems to show a reasonable agreement with the lattice QCD data at higher temperatures. Similarly, the right panel of Fig. 3 shows charmed baryons with strangeness two, and both the IHRG and VDWHRG models give similar values, which underpredict the lQCD estimations. In addition, the EVHRG shows even worse agreement with lQCD. This might be due to the fact that a larger number of charm-strange baryons, with $|S|=2$, are still undiscovered and not predicted by the quark models yet. The similar values of IHRG and VDWHRG point to the fact that the effect of VDW is smaller as the number density of charmed baryons with strangeness two is very low, yet the attractive interactions counter the repulsive interactions reasonably. Future experimental efforts at facilities like LHCb and the upcoming Electron-Ion Collider (EIC) could help constrain these states better.
In Fig. 4 (left panel), we plot the ratio of partial pressure from charmed mesons with strangeness one to total charmed pressure. For the hadron gas case, we have used the total charm pressure as the added pressure from charmed mesons and charmed baryons. However, for the lQCD, they have also added a charm quark pressure, assuming a mixed state near the transition temperature. Due to this contribution from charm quark pressure, the lQCD trend goes down after $T\sim$ 160 MeV. The VDWHRG model explains the lQCD data up to $T\sim$ 160 MeV, but then shows an upward trend after that. This mismatch comes from the neglected charm quark pressure, which seems to dominate after the transition temperature.
Figure 6: Charm susceptibilities ratios, $\kappa_{2}^{C}/\kappa_{1}^{C}$ (left) and $\kappa_{4}^{C}/\kappa_{3}^{C}$ (right), as a function of $\sqrt{s_{\rm NN}}$.
Figure 7: Charm susceptibilities ratios, $\kappa_{3}^{C}/\kappa_{1}^{C}$ (left), $\kappa_{3}^{C}/\kappa_{2}^{C}$ (middle), and $\kappa_{4}^{C}/\kappa_{2}^{C}$ (right), as a function of $\sqrt{s_{\rm NN}}$.
@@
On the other hand, the charm susceptibilities and their ratios as functions of collision energy reveal a multi-faceted picture of QCD matter across different phases. Much like in the baryon sector, in Fig. 6 the increasing trends of $\kappa_{2}^{C}/\kappa_{1}^{C}$ and $\kappa_{4}^{C}/\kappa_{3}^{C}$ with $\sqrt{s_{\rm NN}}$ demonstrate the gradual emergence of correlated charm production mechanisms, transitioning from independent fragmentation at low energies to medium-dominated processes in the QGP phase at high energies. The consistent model hierarchy (VDWHRG << IHRG << EVHRG) highlights how van der Waals interactions most effectively capture the partial thermalization of charm quarks, with attractive forces suppressing fluctuations closer to lQCD predictions. This is complemented by the constancy of $\kappa_{3}^{C}/\kappa_{1}^{C}$ and $\kappa_{4}^{C}/\kappa_{2}^{C}$ saturating beyond center-of-mass energy of 10–30 GeV as shown in the left and right panels of Fig. 7, which serve as a thermodynamic fingerprint of charm’s unique properties, showing that its large mass preserves the charm symmetry and Gaussian statistics despite the phase transition. However, at lower energies, the ratios show an increasing behavior before saturating beyond $\sqrt{s_{\rm{NN}}}=$ 10 GeV due to the consideration of global conservation of charges through the acceptance factor. During this increase, an effect of interaction among the hadrons can be seen in the higher-order cumulant ratio. The decreasing trend of $\kappa_{3}^{C}/\kappa_{2}^{C}$, beyond $\sqrt{s_{\rm{NN}}}=$ 10 GeV, as shown in the middle panel of Fig. 7 and its saturation at high energies reflects the decoupling of charm from baryon number in the $\mu_{B}\to 0$ regime. Here, for center-of-mass energies lower than 10 GeV, an interplay between the effects of global conservation and charm production is observed, which forces the ratio to exhibit an increasing behavior. This comprehensive view positions charm fluctuations as both a sensitive probe of the QCD phase diagram and a stringent test for theoretical models across the entire collision energy spectrum.
@@
The systematic evolution of these ratios with collision energy offers a powerful tool for mapping the QCD phase diagram. Future measurements with improved precision could further constrain the location of the critical point and clarify the temperature dependence of charm hadronization. In addition to this, one key feature that could substantially change these ratios is the charm chemical potential. At RHIC energies, the charm chemical potential ($\mu_{c}$) can be finite due to the nature of charm quark production and conservation in heavy-ion collisions. Charm quarks are predominantly produced in initial hard scatterings rather than through thermal excitation, and their number remains approximately conserved throughout the evolution of the fireball. In thermal models like the HRG, a finite $\mu_{c}$ can be introduced to enforce this conservation, especially when using a grand canonical ensemble. Moreover, local imbalances between charm and anti-charm quarks can develop due to event-by-event fluctuations or limited charm diffusion, further justifying a nonzero $\mu_{c}$ [51]. In thermal fits to charm hadron yields (such as $D^{0}$, $\Lambda_{c}$, or $J/\psi$), incorporating a finite charm chemical potential is necessary to correctly reproduce observed yields under the constraint of fixed net charm. Thus, a thorough look into charm chemical potential and subsequent charm susceptibility estimation is very much timely and necessary.
@@ IV Summary
In this work, we study the thermodynamic properties and charm sector fluctuations of QCD matter using the van der Waals Hadron Resonance Gas (VDWHRG) model. By incorporating both attractive and repulsive interactions between hadrons, the VDWHRG model significantly improves agreement with lattice QCD (lQCD) data, especially in the temperature range up to 180 MeV. We show that the VDWHRG model provides an improved description of charm susceptibilities and partial pressures. Our study includes the estimation of partial pressures of charmed mesons and baryons, both with and without strangeness, and compares them to available lQCD results. While VDWHRG captures the qualitative trends well, especially below $T\sim 160$ MeV, discrepancies at higher temperatures highlight the possible role of deconfined charm degrees of freedom not accounted for in hadron-based models. Additionally, we extend our analysis to baryon-rich regimes, offering insights into charm susceptibilities in regions not easily accessible by lQCD. Furthermore, we study the charm susceptibilities and their ratio as a function of collision energy, which highlights their sensitivity to the underlying QCD phase structure. The observed trend for $\kappa_{2}^{C}/\kappa_{1}^{C}$ and $\kappa_{4}^{C}/\kappa_{3}^{C}$ shows the gradual increase in charm production with an increase in collision energy. Moreover, a saturation behavior in $\kappa_{3}^{C}/\kappa_{2}^{C}$ highlights the decoupling of the charm and baryon number at higher center-of-mass collision energy. For the study of charm susceptibility as a function of center-of-mass energy, we make use of the parametrization established from the light hadron yield. Even though, for the scope of our work, it is sufficient, in the future, one can determine a charm-specific freeze-out condition, for instance, through a statistical hadronization model fit to charm hadron ratios. The upcoming high-luminosity upgrades at the LHC and higher-energy collisions are expected to provide sufficient data for the charm sector to work on such a parametrization.
To summarize, the results presented here provide a strong motivation for further constraining charm-sector interactions through both theoretical and experimental efforts. Theoretically, our predictions for partial pressure at finite $\mu_{B}$ can be tested by future lattice QCD calculations that employ novel methods to circumvent the sign problem. Furthermore, comparing the VDWHRG approach with other models and effective field theories could offer deeper insights into the microscopic origin of these interactions. Experimentally, the most direct path to constrain the VDW parameters is through the measurement of open-charm fluctuations and correlations in heavy-ion collisions, which remains a novel direction and has not been thoroughly explored. Higher-order moments of the distributions of D mesons and $\Lambda_{c}$ baryons are directly sensitive to the charm susceptibilities discussed in this work. Additionally, correlation measurements, such as DpD-p and $D-\Lambda$ femtoscopy, can provide crucial independent constraints on the scattering parameters and effective forces between charm and light hadrons. The upcoming high-statistics data from the LHC Run 3 and the Beam Energy Scan II at RHIC, as well as future programs at FAIR and NICA, will be instrumental in performing these decisive tests, thereby transforming the qualitative success of the VDWHRG model into a quantitative extraction of charm-hadron interactions in the hot and dense medium.